Solve x in the following equation:
4(x+3) - 2(x-5) = 46

Answer by lenny460(1073)

Solve x in the following equation:
4(x+3) - 2(x-5) = 46
4x + 12 - 2x + 10 = 46
2x + 22 = 46
2x = 24
x = 12



lennox obuong
algebra tutor